During an experiment to determine the dimensions of a cylinder, a scholar used a vernier calliper and a balance and recorded the following:Mass of cylinder, m = (52.5 ± 0.1)gHeight of cylinder, h = (18.0 ± 0.1)mmDiameter of cylinder, d = (22.0 ± 0.1)mmDeduce the density of the cylinder, DDeduce the relative error in DDeduce the maximum error in DDeduce the percentage error in m Deduce the percentage error in hDeduce the percentage error in dDeduce the total percentage error in D


1
Expert's answer
2021-01-13T11:37:56-0500

"m=(52.5\pm0.1)g" , "\epsilon=\frac{0.1}{52.5}\cdot100\%=0.2\%"


"h=(18.0\pm0.1)mm" , "\epsilon=\frac{0.1}{18.0}\cdot100\%=0.6\%"


"d=(22.0\pm0.1)mm" , "\epsilon=\frac{0.1}{22.0}\cdot100\%=0.5\%"


"D=\frac{m}{V}=\frac{4m}{\pi d^2h}=\frac{4\cdot0.0525}{3.14\cdot 0.022^2\cdot0.018}=7673(kg/m^3)"


"\frac{\Delta D}{D}=\sqrt{(\frac{\Delta m}{m})^2+(\frac{\Delta h}{h})^2+(2\frac{\Delta d}{d})^2}"


"\frac{\Delta D}{D}=\sqrt{(0.002)^2+(0.006)^2+(2\cdot 0.005)^2}=0.012" or "\epsilon=1.2\%"


"\Delta D=D\cdot0.012=7673\cdot0.012=92(kg/m^3)"


"D=(7670\pm90)kg/m^3" , "\epsilon=1.2\%"